# A General Class of Model-Free Dense Precision Matrix Estimators

### v1

@@ 2 Dimension Reduction
(i). $\min_{1\leq j\leq p}Eigmin(\Sigma_{U,-j})\geq c>0$, where $Eigmin(\Sigma_{U,-j})$ represents the minimum eigenvalue of the matrix $\Sigma_{U,-j}$. Also $\max_{1\leq j\leq p}\sigma_{j}^{2}=\sigma^{2}\leq C<\infty$, and also $\min_{1\leq j\leq p}\sigma_{j}^{2}\geq c>0$.
(ii). Let $\delta_{n}:=\min_{1\leq j\leq p}\|\Sigma_{U,-j}\|_{2}$, and $\delta_{n}$ is nondecreasing in $n$, and also $r_{n}:=\max_{1\leq j\leq p}\|\Sigma_{U,-j}\|_{2}$, where $r_{n}$ is nondecreasing in $n$ but $r_{n}/p\to 0$, and clearly $\delta_{n}\leq r_{n}$.
@@
Assumption 2(i) is also standard in high dimensional statistics literature. It allows for non-singular noise matrices, large noise magnitudes and divergence in the maximum eigenvalue of the (covariance of the) noise. This assumption is very much reflective of financial markets and is used in Caner et al. (2023).
Assumption 3.
(i). Define a positive constant $C_{p}>0$. Then assume the following relation between the number of outcomes, $p$ and the sample size $n$:
$p\leq n^{C_{p}/2}.$
@@
(ii). For each $j=1,\cdots,p$, $A_{-j}:p-1\times K$ has full rank $K$, and $p>K+1$. We impose $\max_{1\leq j\leq p}\|A_{-j}\|_{2}=O(d_{1n})$ with $d_{1n}\to\infty$ as $n\to\infty$, and $d_{1n}/\sqrt{p}\to 0$. Also, $\Sigma_{f}$ has full rank $K$, and $Eigmax(\Sigma_{f})\leq C<\infty$, for C a positive constant.
@@
With $C_{p}>2$, we can obtain $p>n$ and we cover high dimensional cases. In case of $0<C_{p}\leq 2$, we cover low dimensional cases, $p\leq n$. This assumption also allows $p=\gamma n$, with $\gamma\in[0,\infty)$. Full rank $\Sigma_{f}:=E(f_{t}-Ef_{t})(f_{t}-Ef_{t})^{\prime}:K\times K$ matrix is used in Bing et al. (2021), Bunea et al. (2022), reflecting $K<p-1$, where we assume the number of hidden factors to be less than number of assets minus one. We allow the spectral norm of $A_{-j}:p-1\times K$ matrix to be diverging in $n$, reflective of the high dimensional nature of $A_{-j}$, where both dimensions are allowed to grow.
Assumption 4.
@@
(i). For each $j=1,\cdots,p$ let $u_{jt}$ be zero mean $\sigma_{j}\gamma_{e}$ subgaussian random variable which means for each $j=1,\cdots,p$ $Eexp(t_{1}u_{jt})\leq exp(t_{1}^{2}\sigma_{j}^{2}\gamma_{e}^{2}/2)$ for all $t_{1}\in R$. Define the zero mean vector $U_{-j,t}:=\Sigma_{U,-j}^{1/2}\tilde{U}_{-j,t}$, where $\tilde{U}_{-j,t}$ ($p-1\times 1$ vector) is $\gamma_{wj}$ sub-Gaussian vector which means for any unit vector ($p-1$ dimension) $v$, $v^{\prime}\tilde{U}_{-j,t}$ is $\gamma_{wj}$ sub-Gaussian. Also we assume $E\tilde{U}_{-j,t}\tilde{U}_{-j,t}^{\prime}=I_{p-1}$. Without losing any generality in the proofs we can use $\gamma_{w}$ constant instead of changing constants for each $j$, and $\gamma_{wj}$.
(ii). Define $f_{t}:=\Sigma_{f}^{1/2}\tilde{f}_{t}$, where $\tilde{f}_{t}:K\times 1$, is a subgaussian vector with $E[\tilde{f}_{t}\tilde{f}_{t}^{\prime}]=I_{K}$, and $\max_{1\leq k\leq K}|Ef_{tk}|\leq C<\infty$, where $C>0$ is a positive constant.
(iii). We assume $\|A\|_{\infty}\leq C<\infty$.
@@
Assumption 4 is standard assumption used in Bing et al. (2021) Bunea et al. (2022). This assumption is needed for quadratic form high dimensional probability inequalities. It should be noted that we also relax the non-zero factor mean assumption used in earlier literature in Bing et al. (2021), and Bunea et al. (2022).
Define the uniform version of the signal to noise ratio
@@
with $\lambda_{K}(A_{-j}\Sigma_{f}A_{-j}^{\prime})$ representing the $K$ th largest eigenvalue of $A_{-j}\Sigma_{f}A_{-j}^{\prime}:p-1\times p-1$ matrix, and that is the smallest positive eigenvalue of that matrix since $p-1>K$, and $\Sigma_{f}:K\times K$ matrix.
We have a discussion about the role of the signal-to-noise ratio in a specific scenario. Since $A_{-j}:p-1\times K$ for all matrices $j=1,\cdots,p$, via Assumption 3, if the largest eigenvalue of $A_{-j}^{\prime}A_{-j}$ grows at the same rate $K$-th largest eigenvalue, we can assume $\lambda_{K}(A_{-j}^{\prime}A_{-j})\asymp d_{1n}^{2}$ and since we impose $max_{1\leq j\leq p}\|\Sigma_{U,-j}\|_{2}=r_{n}$ by Assumption 3 then
@@
$\bar{\xi}\geq\frac{cd_{1n}^{2}}{r_{n}}.$ (8)
Depending on how fast/slow $d_{1n}$, and $r_{n}$ grow the signal-to-noise ratio lower bound in (8) can fluctuate between fairly small and fairly large quantities.
Assumption 5 below covers diagonal estimation of precision matrix. Note that $\bar{r},\bar{\Psi},\bar{\eta}$ are three random sequences that are explained in detail before Theorem 1, and they represent the complexity, prediction error, and prediction performance of the used technique respectively. In special cases of RRE, and adaptive PCR, we will show non-random upper bounds for these three key quantities.
To state Assumption 5 below, we find it convenient to introduce two rates, $r_{w1}$, $r_{w2}$ where $r_{w2}$ is the square root of the rate for off-diagonal terms in the precision matrix estimation. The rate $r_{w1}$ (compared to $r_{w2}$) is also used in estimation for the off-diagonal terms.
$r_{w1}:=\max\left(\sqrt{\frac{logn+\bar{r}}{n\bar{\eta}}},\sqrt{\frac{K}{\bar{\xi}\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}}\frac{\bar{\Psi}}{\bar{\eta}}\frac{1}{\delta_{n}}},\sqrt{\frac{K}{\bar{\xi}}\frac{1}{\bar{\eta}}}\right).$ (9)
$r_{w2}:=\max\left(r_{w1},\sqrt{\frac{K}{\bar{\xi}}}\right).$ (10)
Note the only difference between $r_{w1},r_{w2}$ is the second term in parentheses in (9)(10), since $\delta_{n}$ is nondecreasing in $n$, we can tell that $r_{w2}$ is the larger rate.
@@
Assumption 5.
(i). $\frac{d_{1n}}{\sqrt{p}}\frac{1}{\sqrt{K}}\max(1,\frac{\sqrt{logn}}{\sqrt{logp}})\to 0$.
(ii). $\sqrt{\frac{K^{3}}{n}}\to 0,\quad K\frac{\sqrt{max(logp,logn)}}{n^{1/2}}\to 0.$
(iii). $K\sqrt{\frac{p(logp)}{n}}r_{w1}\stackrel{{\scriptstyle p}}{{\to}}0$, and $d_{2n}r_{w2}\stackrel{{\scriptstyle p}}{{\to}}0$, and $d_{2n}$ is nondecreasing in $n$.
While the meaning of Assumption 5 will become clearer when we get to our final results, a few introductory remarks regarding the origin of the assumption are in place. First, Assumption 5(i) is a technical assumption that facilitates the proof of Theorem 2 for the finite sample results. [^4] One should note that this is a strengthened version of Assumption 3 when p and K are small; with larger p compared to n, Assumption 3 will suffice and this part of Assumption 5 will not be needed. Assumption 5(ii) relates to the tradeoff between number of factors and time span. In particular, it points out that the number of factors should be much smaller than the time span. Assumption 5(iii) involves stochastic terms and relates to consistency and rate of convergence of the off-diagonal terms estimates. As we will see later on, the role of Assumption 5(iii) will be discussed separately for adaptive PCR and RRE estimators.
@@ 3 Linear General Class of Estimators
We propose estimating the rows of the precision matrix of asset returns with our estimator. To show consistency of our estimators in high dimensional context we need to start with a precision matrix formula that can be tied to hidden factors. This is not a trivial task. One of the main obstacles is establishing relation between the matrix inversion and linear regression within hidden factor context. The difficulty is that formula for main diagonal terms and off-diagonal terms of the precision matrix is different and this type of approach tying hidden factors and general linear class of estimators with precision matrix estimation has not been done before. This is one of the contributions of our paper. Define the following parameter vector which will form a key term in precision matrix formula:
@@
This definition shows that we can relate to linear regression context. Note that, the definition of $\alpha_{j}^{*}$ is not coming from the hidden factor model directly in (2)(3). We also need the following definitions before the next Lemma. Let $\Sigma_{f}=\Sigma_{f}^{1/2}\Sigma_{f}^{1/2}$ since $\Sigma_{f}$ is a positive definite matrix. Set $\bar{A}_{-j}:=A_{-j}\Sigma_{f}^{1/2}$, and $\bar{a}_{j}:=\Sigma_{f}^{1/2}a_{j}$, and $\bar{G}_{j}:=I_{K}+\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}.$ Next, we show how they relate to hidden factor structure. Let $\Theta_{j,j}$ represent the main $j$ th diagonal term, and $\Theta_{j,-j}$ represent the $j$ th row without $j$ th element.
Lemma 1.
@@
Remark 1.
We make the following observations about the precision matrix formula and its relation to hidden factors:
  • One key result is that we will not impose arbitrary zero-exact sparsity restrictions on $\alpha_{j}^{*}$ as done in nodewise regression in Meinshausen and Bühlmann (2006), Callot et al. (2021).
  • Note that our precision matrix formula is not a direct extension nor an alternative representation of Sherman-Morrison-Woodbury matrix inversion formula. Rather, we start from a matrix algebra based formula and then impose hidden factor structure for outcomes to derive the precision matrix. The hidden factor structure is needed when we derive consistency of our estimator tied to the precision matrix formula.
  • Note that the formulas for $\alpha_{j}^{*}$ and $\tau_{j}^{2}$ are expressed in terms of the hidden factor based coefficients. This represents a novel contribution compared to existing literature—such as Callot et al. (2021), Caner et al. (2023), and Caner and Kock (2018)—which analyze settings without factors, consider errors in an observed factor model, or focus on regressors in a standard linear regression framework, respectively.
Connecting further to the existing literature, one should note that Bing et al. (2021) propose linear regression coefficient estimation and analyze the risk of an asset in a hidden factor model. However, neither the precision matrix estimation nor the above formulae are discussed therein. The main precision matrix estimation difficulty is that there are p rows which are tied to hidden factors. This leaves as widely open questions related to how and why linear regression estimators typically employed in hidden factor models should generalize to the precision matrix estimation. Lemma 1 and Theorem 1 below show how we manage to find a way to base the estimation on hidden factors but without actually imposing any identification conditions on hidden factors.
In that respect for each $j=1,\cdots,p$, we define
@@
where $(.)^{+}$ is the Moore-Penrose inverse of a matrix, and $\hat{B}:p-1\times q$ with $q\leq p-1$ matrix, and the second equality is by (134) of Bunea et al. (2022). Dimension $q$ will be tied to specific estimators, and immediately below, we give examples about that. We let $\hat{B}$ be a deterministic matrix or a matrix that depends on $Y_{-j}$. Our estimator is based on analysis of precision matrix in Lemma 1. With different values of $\hat{B}$ we obtain different estimators. To see that, $\hat{B}=I_{p-1}$ is the ridgeless regression estimator (RRE) which is analyzed also after this section separately, so $q=p-1$. In case of adaptive PCR, $\hat{B}=V_{1,k_{j}}$ with $V_{1,k_{j}}$ being the $p-1\times k_{j}$ matrix where the columns are the eigenvectors of $Y_{-j}^{\prime}Y_{-j}/n$ matrix corresponding to the largest $k_{j}$ eigenvalues, and $q=k_{j}$, $j=1,\cdots,p$. We do not subscript $\hat{B}$ with $j$ to simplify the notation. Let us define the reciprocal of the main diagonal terms in the precision matrix based on (B.11) [^5] Note that only in case of RRE estimator as the subcase when $p>n$ the estimator simplifies and we benefit from that expression.

### v2

@@ 2 Dimension Reduction
(i). $\min_{1\leq j\leq p}Eigmin(\Sigma_{U,-j})\geq c>0$, where $Eigmin(\Sigma_{U,-j})$ represents the minimum eigenvalue of the matrix $\Sigma_{U,-j}$. Also $\max_{1\leq j\leq p}\sigma_{j}^{2}=\sigma^{2}<\Gamma<\infty$, with $\Gamma$ as a large positive constant, and also $\min_{1\leq j\leq p}\sigma_{j}^{2}\geq c>0$. (ii). Let $\beta:=\min_{1\leq j\leq p}\|\Sigma_{U,-j}\|_{2}$, and $\beta>0$ is a large positive constant , and also $\max_{1\leq j\leq p}\|\Sigma_{U,-j}\|_{2}<\Gamma$, where $\Gamma>0$ is described in (i). Clearly $\beta<\Gamma$.
@@
Assumption 2(i) is also standard in high dimensional statistics literature. It allows for non-singular noise matrices, large noise magnitudes. Since $\Sigma_{U,-j}$ is diagonal and diagonals have error variances, we use a common upper bound for a single variance in (i) and maximum variances over errors in (ii). This large positive $\Gamma$ is very much reflective of financial markets.
Assumption 3
@@
(i). Define a positive constant $C_{p}>0$. Then assume the following relation between the number of outcomes, $p$ and the sample size $n$: $p\leq n^{C_{p}/2}.$ (ii). For each $j=1,\cdots,p$, $A_{-j}:p-1\times K$ has full rank $K$, and $p>K+1$. We impose $\max_{1\leq j\leq p}\|A_{-j}\|_{2}=O(\sqrt{p})$. Also, $\Sigma_{f}$ has full rank $K$, and $Eigmax(\Sigma_{f})\leq C<\infty$, for C a positive constant.
@@
With $C_{p}>2$, we can obtain $p>n$ and we cover high dimensional cases. In case of $0<C_{p}\leq 2$, we cover low dimensional cases, $p\leq n$. This assumption also allows $p=\gamma n$, with $\gamma\in[0,\infty)$. We allow the spectral norm of $A_{-j}:p-1\times K$ matrix to be diverging in $n$, reflective of the high dimensional nature of $A_{-j}$, where both dimensions are allowed to grow. Specifically in (ii), we allow for pervasive factors with $\max_{j}\|A_{-j}\|_{2}=C\sqrt{p}$, $C$ being a generic positive constant as well as weak factor structures if $\max_{j}\|A_{-j}\|_{2}=O(d_{1n})$, with $d_{1n}/\sqrt{p}\to 0$ with $p\to\infty,n\to\infty$. Hence our assumption is a weak one covering several possibilities.
Assumption 4
@@
(i). For each $j=1,\cdots,p$ let $u_{jt}$ be zero mean $\sigma_{j}\gamma_{e}$ subgaussian random variable which means for each $j=1,\cdots,p$ $Eexp(t_{1}u_{jt})\leq exp(t_{1}^{2}\sigma_{j}^{2}\gamma_{e}^{2}/2)$ for all $t_{1}\in R$. Define the zero mean vector $U_{-j,t}:=\Sigma_{U,-j}^{1/2}\tilde{U}_{-j,t}$, where $\tilde{U}_{-j,t}$ ($p-1\times 1$ vector) is $\gamma_{wj}$ sub-Gaussian vector which means for any unit vector (p1p-1 dimension) $v$, $v^{\prime}\tilde{U}_{-j,t}$ is $\gamma_{wj}$ sub-Gaussian. Also we assume $E\tilde{U}_{-j,t}\tilde{U}_{-j,t}^{\prime}=I_{p-1}$. Without losing any generality in the proofs we can use $\gamma_{w}$ constant instead of changing constants for each $j$, and $\gamma_{wj}$. (ii). Define $f_{t}:=\Sigma_{f}^{1/2}\tilde{f}_{t}$, where $\tilde{f}_{t}:K\times 1$, is a subgaussian vector with $E[\tilde{f}_{t}\tilde{f}_{t}^{\prime}]=I_{K}$, and $\max_{1\leq k\leq K}|Ef_{tk}|\leq C<\infty$, where $C>0$ is a positive constant. (iii). We assume $\|A\|_{\infty}\leq C<\infty$.
@@
Assumption 4 is standard assumption used in Bing et al. (2021) Bunea et al. (2022). This assumption is needed for quadratic form high dimensional probability inequalities. It should be noted that we also relax the non-zero factor mean assumption used in earlier literature in Bing et al. (2021), and Bunea et al. (2022). Define the uniform version of the signal to noise ratio
@@
with $\lambda_{K}(A_{-j}\Sigma_{f}A_{-j}^{\prime})$ representing the $K$ th largest eigenvalue of $A_{-j}\Sigma_{f}A_{-j}^{\prime}:p-1\times p-1$ matrix, and that is the smallest positive eigenvalue of that matrix since p1>Kp-1>K, and $\Sigma_{f}:K\times K$ matrix. We have a discussion about the role of the signal-to-noise ratio in a specific scenario. Let us assume a weaker factor structure in Assumption 3(ii), which is $\max_{j}\|A_{-j}\|_{2}=O(d_{1n})$. Since $A_{-j}:p-1\times K$ for all matrices $j=1,\cdots,p$, via Assumption 3, if the largest eigenvalue of $A_{-j}^{\prime}A_{-j}$ grows at the same rate $K$-th largest eigenvalue, we can assume $\lambda_{K}(A_{-j}^{\prime}A_{-j})\asymp d_{1n}^{2}$ and since we impose $max_{1\leq j\leq p}\|\Sigma_{U,-j}\|_{2}<\Gamma$ by Assumption 3 then
@@
$\bar{\xi}\geq\frac{cd_{1n}^{2}}{\Gamma}.$ (8)
Depending on how fast/slow $d_{1n}$ grows the signal-to-noise ratio lower bound in (8) can fluctuate between fairly small and fairly large quantities. Assumption 5 below covers diagonal estimation of precision matrix. Note that $\bar{r},\bar{\Psi},\bar{\eta}$ are three random sequences that are explained in detail before Theorem 1, and they represent the complexity, bias, and size of the signal respectively. In special cases of RRE, and adaptive PCR, we will show non-random upper bounds for these three key quantities. To state Assumption 5 below, we find it convenient to introduce rate of convergence, $r_{w1}$ which is used for off-diagonal term estimation.
$r_{w1}:=\max\left(\sqrt{\frac{logn+\bar{r}}{n\bar{\eta}}},\sqrt{\frac{K}{\bar{\xi}}},\sqrt{\frac{K}{\bar{\xi}}\frac{\bar{\Psi}}{\bar{\eta}}},\sqrt{\frac{K}{\bar{\xi}}\frac{1}{\bar{\eta}}}\right).$ (9)
@@
Assumption 5
(i). $\frac{1}{\sqrt{K}}\frac{\sqrt{logn}}{\sqrt{logp}}\to 0$. (ii). $\sqrt{\frac{K^{3}}{n}}\to 0,\quad K\frac{\sqrt{max(logp,logn)}}{n^{1/2}}\to 0.$ (iii). $K\sqrt{\frac{p(logp)}{n}}r_{w1}\stackrel{{\scriptstyle p}}{{\to}}0$, and $d_{2n}r_{w1}\stackrel{{\scriptstyle p}}{{\to}}0$, and $d_{2n}$ is nondecreasing in $n$.
While the meaning of Assumption 5 will become clearer when we get to our final results, a few introductory remarks regarding the origin of the assumption are in place. First, Assumption 5(i) is a technical assumption that facilitates the proof of Theorem 2 for the finite sample results. Assumption 5(ii) relates to the tradeoff between number of factors and time span. In particular, it points out that the number of factors should be much smaller than the time span. Assumption 5(iii) involves stochastic terms and relates to consistency and rate of convergence of the off-diagonal terms estimates. As we will see later on, the role of Assumption 5(iii) will be discussed separately for adaptive PCR and RRE estimators. But to provide some basic intuition in a simple subcase, let $K=1$, and $\|\Sigma_{U,-j}\|_{2}=I_{p-1}$, with $var(f_{t})=1$, and set $r_{w1}:=1/\sqrt{\xi}$
$d_{2n}r_{w1}=\frac{\max_{j}\sqrt{\sum_{i\neq j}(A_{i,-j}a_{j})^{2}}}{\min_{j}\sqrt{\sum_{i\neq j}A_{i,-j}^{2}}}\to 0,$
where $A_{i,-j}$ is $i$ th element of vector $A_{-j}$ (see that this is a vector since $K=1$). So this is basically factor loadings correlations among themselves declining to zero.
@@ 3 Linear General Class of Estimators
We propose estimating the rows of the precision matrix of asset returns with our estimator. To show consistency of our estimators in high dimensional context we need to start with a precision matrix formula that can be tied to hidden factors. One of the main obstacles is establishing relation between the matrix inversion and linear regression within hidden factor context. The difficulty is that formula for main diagonal terms and off-diagonal terms of the precision matrix is different and this type of approach tying hidden factors and general linear class of estimators with precision matrix estimation has not been done before. Define the following parameter vector which will form a key term in precision matrix formula:
@@
This definition shows that we can relate to linear regression context. Note that, the definition of $\alpha_{j}^{*}$ is not coming from the hidden factor model directly in (2)(3). We also need the following definitions before the next Lemma. Let $\Sigma_{f}=\Sigma_{f}^{1/2}\Sigma_{f}^{1/2}$ since $\Sigma_{f}$ is a positive definite matrix. Set $\bar{A}_{-j}:=A_{-j}\Sigma_{f}^{1/2}$, and $\bar{a}_{j}:=\Sigma_{f}^{1/2}a_{j}$, and $\bar{G}_{j}:=I_{K}+\bar{A}_{-j}^{\prime}\Sigma_{U,-j}^{-1}\bar{A}_{-j}.$ Next, we show how they relate to hidden factor structure. Let $\Theta_{j,j}$ represent the main $j$ th diagonal term, and $\Theta_{j,-j}$ represent the $j$ th row without $j$ th element. One key result in Lemma 1 is that we will not impose arbitrary zero-exact sparsity restrictions on $\alpha_{j}^{*}$ as done in nodewise regression in Meinshausen and Bühlmann (2006), Callot et al. (2021).
Lemma 1
@@
Connecting further to the existing literature, one should note that Bing et al. (2021) propose linear regression coefficient estimation and analyze the risk of an asset in a hidden factor model. However, neither the precision matrix estimation nor the above formulae are discussed therein. Lemma 1 and Theorem 1 below show how we manage to find a way to base the estimation on hidden factors but without actually imposing any identification conditions on hidden factors. In that respect for each $j=1,\cdots,p$, we define
@@
where $(.)^{+}$ is the Moore-Penrose inverse of a matrix, and $\hat{B}:p-1\times q$ with $q\leq p-1$ matrix, and the second equality is by (134) of Bunea et al. (2022). Dimension $q$ will be tied to specific estimators, and immediately below, we give examples about that. We let $\hat{B}$ be a deterministic matrix or a matrix that depends on $Y_{-j}$. Our estimator is based on analysis of precision matrix in Lemma 1.
Note that one of the main contributions of this paper is to link estimators (10) with $\alpha_{j}^{*}$ which is related to the matrix inversion. With different values of $\hat{B}$ we obtain different estimators. To see that, $\hat{B}=I_{p-1}$ is the ridgeless regression estimator (RRE) which is analyzed also after this section separately, so q=p1q=p-1. In case of adaptive PCR, $\hat{B}=V_{1,k_{j}}$ with $V_{1,k_{j}}$ being the $p-1\times k_{j}$ matrix where the columns are the eigenvectors of $Y_{-j}^{\prime}Y_{-j}/n$ matrix corresponding to the largest $k_{j}$ eigenvalues, and $q=k_{j}$, $j=1,\cdots,p$. We do not subscript $\hat{B}$ with $j$ to simplify the notation.  [^2] An optimal $\hat{B}$ may be analyzed but without knowing its structure it may be difficult and it is beyond current scope of the paper. Let us define the reciprocal of the main diagonal terms in the precision matrix based on (B.11) [^3] Note that only in case of RRE estimator as the subcase when p1>np-1>n the estimator simplifies and we benefit from that expression.